what is the answer of 3x+2y=-3 1x+1/3y=-4 by substitution
Answer by Cromlix(4381)

3x + 2y = -3
.............
x + 1/3y = -4
x = -4 - 1/3y
Substitute into
3x + 2y = -3
3(-4 - 1/3y) + 2y = -3
-12 - y + 2y = -3
Collect like terms:
- y + 2y = - 3 + 12
y = 9
Substitute y = 9 into
x + 1/3y = -4
x + 1/3(9) = -4
x + 3 = - 4
x = -4 -3
x = -7
{-7,9}
Hope this helps.
:-)